OpenStudy (addonexus):

Two angles are supplementary. One is 48 degrees more than 3 times the other. find the measures of the angles. *supplementary angles are angles whose sum equals 180 degrees*

OpenStudy (anonymous):

do you have choices ?

OpenStudy (addonexus):

no, it's fill-in.

OpenStudy (anonymous):

alright il help

OpenStudy (anonymous):

you have to do 180-48 which is 132

OpenStudy (addonexus):

Ok

OpenStudy (anonymous):

for example let x be the smaller angle and 3x be the bigger angle write the equation 3x-x = 180 combine like terms 4x=180 and divide both sidesby 4 and you get 45 smaller angle would be 45 degrees and the bigger angle would be 135 45+135= 180 does that help you understand

OpenStudy (addonexus):

135 -48 = 87 and 87/3 = 29 The bigger angle has to be 48 more than 3 times the smaller angle. so if it was 135 for the bigger angle, 29 would be the smaller angle and 135 + 29 does not equal 180

OpenStudy (anonymous):

you cant subtract it always add

OpenStudy (anonymous):

nevermind i see what you did

OpenStudy (anonymous):

guys and ladys can you sighn up for me this is a present for helping me erlier its free stuff and im not kiding https://www.listia.com/signup/7448912

OpenStudy (anonymous):

will you?